### v1

# Synthetic multi-inversion time magnetic resonance images for visualization of subcortical structures

Abstract
Purpose: Visualization of subcortical gray matter is essential in neuroscience and clinical practice, particularly for disease understanding and surgical planning. While multi-inversion time (multi-TI) T1-weighted (T1-w) magnetic resonance (MR) imaging improves visualization, it is rarely acquired in clinical settings.
Approach: We present SyMTIC (Synthetic Multi-TI Contrasts), a deep learning method that generates synthetic multi-TI images using routinely acquired T1-w, T2-weighted (T2-w), and FLAIR images. Our approach combines image translation via deep neural networks with imaging physics to estimate longitudinal relaxation time (T1) and proton density (PD) maps. These maps are then used to compute multi-TI images with arbitrary inversion times.
Results: SyMTIC was trained using paired MPRAGE and FGATIR images along with T2-w and FLAIR images. It accurately synthesized multi-TI images from standard clinical inputs, achieving image quality comparable to that from explicitly acquired multi-TI data. The synthetic images, especially for TI values between 400–800 ms, enhanced visualization of subcortical structures and improved segmentation of thalamic nuclei.
Conclusion: SyMTIC enables robust generation of high-quality multi-TI images from routine MR contrasts. It generalizes well to varied clinical datasets, including those with missing FLAIR images or unknown parameters, offering a practical solution for improving brain MR image visualization and analysis.
keywords:
MRI, image synthesis, brain
[…]
1 Introduction
Visualization of subcortical structures using magnetic resonance imaging is important for understanding the causes and progression of various diseases [1] and for planning and assessing the efficacy of their treatment [2]. Most imaging protocols used to image the thalamus, its nuclei, the basal ganglia, and related cortical networks require special pulse sequences such as quantitative susceptibility mapping (QSM) [3] and fast gray matter acquisition T1 inversion recovery (FGATIR) [4] or specialized hardware such as 7T scanners [5]. The most common approach to thalamic nuclei delineation relies on nonlinear deformation of atlases to the acquired MR images, a process which is laden with high variability [6]. White matter nulled (WMn) T1-weighted (T1-w) imaging, such as the FGATIR image shown in Fig. 1(d), is often used in surgical planning for deep brain stimulation (DBS) [2] and is employed in some automatic methods for segmentation of thalamic nuclei [7, 8]. However, FGATIR images are not commonly acquired in routine clinical scans despite providing better contrast than conventional images. In this paper, we present a new way to visualize and delineate thalamic nuclei and related structures using conventional clinical magnetic resonance (MR) imaging.
Figure 1: Resulting images from a 3D MPRAGE sequence with two different inversion times (TIs). (a) The cerebrospinal fluid nulled magnetization prepared rapid acquisition with gradient echo sequence (MPRAGE) which is the same sequence with a TI of 1,200ms and (d) a fast gray matter acquisition T1 inversion recovery (FGATIR) sequence with a TI of 400ms. Shown in (b) and (c) are zoomed versions of (a) and (b), respectively, that are focused on the thalamus. The conventional long TI, MPRAGE image shows relatively poor contrast between subthalamic nuclei, whereas the short TI, FGATIR images have sufficient contrast to visualize and segment these structures.
The 3D magnetization prepared rapid acquisition with gradient echo (MPRAGE) image [9] is routinely acquired in neuroimaging, largely because of its excellent gray matter (GM)/white matter (WM) contrast as well as its fast acquisition time and high resolution. It is conventionally acquired to yield a T1-w contrast using a $180^{\circ}$ inversion pulse and an inversion time (TI) that will null the signal from cerebrospinal fluid, as shown in Fig. 1(a). The FGATIR pulse sequence is a 3D MPRAGE sequence with a much shorter TI (typically around 400 ms at 3T), which is selected to approximately null the WM signal. Although a conventional MPRAGE image shows relatively poor contrast of the subthalamic nuclei, striatum, globus pallidus interna, nucleus accumbens, and internal capsule, FGATIR images have been shown to have sufficient contrast to target and/or segment these critical brain structures [4].
Despite the utility of FGATIR and similar sequences in improving subcortical structure contrast, these specialized imaging protocols are not routinely included in clinical workflows due to their additional acquisition time, limited availability in standard imaging protocols, and challenges in consistent implementation across diverse clinical sites. Consequently, there is a pressing need for methods that can enhance the visualization and delineation of subcortical structures using only routinely acquired MR sequences.
The difference between the MPRAGE and FGATIR sequences comes from their different TI values, which lead to significant differences in image contrast. Acquiring additional MPRAGE images with different TI values is typically not feasible in clinical practice due to time constraints. However, when at least two MPRAGE images with different TIs (e.g., conventional MPRAGE and FGATIR) are available, they can be used to estimate T1 and PD maps. These tissue parameter maps can then be used with imaging equations to calculate contrast-weighted MR images with arbitrary parameters [10, 11] and multi-TI images. Jog et al. [12] introduced an approach combining machine learning-based synthesis with imaging equations, where estimated pulse sequence parameters were used to modify an atlas image to improve synthesis quality. However, their method relied on prior tissue segmentation to estimate these parameters, making it less applicable in clinical settings where automated segmentation is not always reliable. Unlike their approach, our method directly estimates T1 and PD maps from commonly acquired images without requiring segmentation.
Some recent work [13, 14, 15] has on synthesizing FGATIR images without segmentation. Umapathy et al. [13] used a single MPRAGE image to synthesize an FGATIR image using a deep network, showing better performance in thalamus segmentation on the synthesized FGATIR than on the acquired MPRAGE. Moya-Sáez et al. [14] calculated T1, T2, and PD parameter maps from T1-w and T2-w images, training a deep learning approach with synthetic data to generate an FGATIR image. They observed a degradation in performance of their model when using non skull-stripped images. Tohidi et al. [15] introduced an additional loss during network training, using an imaging equation when synthesizing the T1 and PD parameter maps, reporting superior performance despite using a small dataset and skull-stripped images. Skull-stripping has previously been reported to make the synthesis task easier [16]; however, it eliminates skull adjacent regions of interest including the subarachnoid space. Our previous work [17] synthesized non-skull-stripped FGATIR images using supervised and unsupervised approaches, with a 3D U-Net showing better performance than using multi-atlas registration [18]. However, this approach did not solve for quantitative parameter maps and was limited to directly synthesizing an FGATIR image. More closely related to this work is Middlebrooks et al. [19], which targeted surgical planning for DBS. They calculated T1 maps from the MP2RAGE sequence, which provides two images with different TIs, and used a simplified imaging equation, which nullified the PD component. As the approach of Middlebrooks et al. requires a MP2RAGE sequence, it has limited applicability as such data is not routinely acquired nor is it commonly available in public datasets.
Building on our previous work [15], we present SyMTIC (Synthetic Multi-TI Contrasts), a method that synthesizes T1 and PD parameter maps from three commonly acquired MR images: T1-w, T2-weighted (T2-w), and fluid-attenuated inversion recovery (FLAIR). Unlike prior approaches, SyMTIC does not rely on specialized acquisitions or skull-stripped data, and it generalizes across domains using harmonization. The key contributions of this work are:
  • We introduce a physics-informed learning framework that enables synthesis of multi-TI images at arbitrary TIs, rather than at a fixed set of values.
  • We demonstrate three different training strategies that optimize parameter map accuracy and image synthesis quality.
  • We retain the full head, including the skull region, to support downstream analyses such as skull-based registration or MRgFUS planning, which most existing methods exclude.
  • By incorporating contrast imputation through harmonization, we show that SyMTIC remains effective even when one of the key input images (FLAIR) is missing.
  • We evaluate SyMTIC on out-of-domain datasets, demonstrating that the model generalizes well to clinical images acquired with different scanners and protocols.
Our model is open source and publicly available from https://github.com/UponAcceptance.
2 Methods
2.1 Overview
Figure 2 provides an overview of SyMTIC, which can be summarized as consisting of four steps: harmonization and imputation of out-of-domain images using HACA3 [20] (Fig. 2A), in-domain images input to the U-Net model (Fig. 2B), synthesis of the T1 and PD parameter maps (Fig. 2C), and application of imaging equations (Fig. 2D) to generate multi-TI images. We note that this framework pertains solely to the inference phase, and a distinct setup is employed during the training phase to estimate the parameter maps. Harmonization and imputation are only used when the images are out-of-domain or there is a missing image contrast (e.g., no acquired FLAIR). In the following sections, we discuss our data, preprocessing steps, imaging equations, synthesis model training, and test time harmonization and imputation.
Figure 2: Overview of SyMTIC. (A) Harmonization and imputation of out-of-domain images using HACA3 (if needed). (B) In-domain images input to the U-Net model. (C) Synthesis of the T1 and PD parameter maps. (D) Calculation of multi-TI images using the imaging equation and specific TIs.
2.2 Data
Our training dataset consists of four brain MR image contrasts that have been acquired for 23 subjects. These contrasts include MPRAGE, FGATIR, T2-w, and FLAIR. Example images of these four contrasts are shown in Fig. 3. The MPRAGE and FGATIR images are acquired with the same imaging parameters (TR = 4,000ms, echo time = 3.37ms, flip angle = $6^{\circ}$, and magnetic field strength = 3 Telsa) with the exception of their inversion time (TI), which are TI = 1,400ms for the MPRAGE and TI = 400ms for the FGATIR. For training the networks described in Sec. 2.5, we used a 5-fold cross validation. We divided our 23 subjects into 14 training subjects, 4 validation subjects, and 5 testing subjects. For the 5th-fold, only 3 testing subjects were used to allow each subject to be tested once. The validation set was used to monitor the loss during training and for early stopping.
MPRAGE FGATIR T2-w FLAIR
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: Example images from the same subject, from left to right, are the MPRAGE, FGATIR, T2-w, and FLAIR.
2.3 Preprocessing
Convolutional neural networks (CNNs) generally perform better when bias field correction and intensity normalization are applied, as these preprocessing steps help mitigate image intensity variations that are unrelated to tissue properties. We use N4 bias field correction [21], which is a robust and widely used method, followed by rigid registration to the MNI152 atlas using ANTs [22]. Intensity normalization was performed using white matter mean normalization as described in Reinhold et al. [23], following the approach of Tohidi et al. [15]. For our methodology, preprocessing must be handled carefully for the MPRAGE and FGATIR images due to their direct mathematical relationship. Since these images are acquired with identical imaging parameters except for their TIs, it is essential to maintain their quantitative relationship to ensure that the imaging equations remain valid. To perform bias field correction while preserving this relationship, we first estimate the bias field for both the MPRAGE and FGATIR images separately and then compute the geometric mean of these two bias fields. This single bias field is then applied to correct each image, ensuring that the relative intensity scaling between them remains consistent. For white matter mean normalization, we first normalize the MPRAGE image and then apply the same normalization factor to the FGATIR image, maintaining their relative intensity relationship.
2.4 Physics-Based Synthesis of Multi-TI Images
As described in Hornak [24], the intensity $I$ at any voxel $v$ in an inversion recovery sequence can be modeled by the following equation:
$I(v)=M_{PD}(v)\left[1-2\exp{\left(\frac{-T_{I}}{M_{T1}(v)}\right)}+\exp{\left(\frac{-T_{R}}{M_{T1}(v)}\right)}\right],$ (1)
where $I(v)$ is the image intensity at $v$, $M_{PD}(v)$ is the PD map value at $v$, $M_{T1}(v)$ is the T1 map value at $v$, and $T_{R}$ is the repetition time, which is 4,000ms for both our MPRAGE and FGATIR images. Since the repetition time is fixed and known from the acquisition protocol, and the TI values for the acquired MPRAGE $T_{I1}$ and FGATIR $T_{I2}$ images are also known, we can solve for the two unknowns, $M_{PD}$ and $M_{T1}$ at each voxel. We rewrite Equation 1 as:
$f(x,y;a,b)=x\left[1-2\exp{\left(\frac{-a\phantom{-}}{y}\right)}+\exp{\left(\frac{-b\phantom{-}}{y}\right)}\right],$ (2)
where $x$ and $y$ correspond to $M_{PD}$ and $M_{T_{1}}$, respectively. We then solve the following optimization problem at each voxel $v$ using a least squares formulation:
$(M_{PD}(v),M_{T1}(v))$ $=$ $\operatorname*{arg\,min}_{(x,y)\in\mathbb{R}^{2}}\left(\left(I_{1}(v)-f(x,y;T_{I1},T_{R})\right)^{2}\right.$ (3)
$\left.\qquad+\left(I_{2}(v)-f(x,y;T_{I2},T_{R})\right)^{2}\right),$
where $I_{1}$ is the acquired MPRAGE image, $I_{2}$ is the acquired FGATIR image, and $T_{I1}$ and $T_{I2}$ are their corresponding inversion times. Although Eq. 3 assumes signed signal values, most MR images are reconstructed and stored as magnitude images. In particular, the FGATIR sequence is acquired at a very short TI (e.g., TI = 400 ms), where the longitudinal magnetization is still negative for many tissues prior to readout. To correctly solve for ${M_{PD}}$ and ${M_{T_{1}}}$ using Eq. 3, we negate the FGATIR image values prior to optimization to reflect the signed nature of the signal predicted by the inversion recovery equation. This detail is critical for accurate parameter estimation; however, for rendering synthetic multi-TI images from Eq. 1, we apply the absolute value to match the magnitude format of clinically acquired MR images.
After computing $M_{PD}$ and $M_{T1}$ for all voxels, we can use Eq. 1 to calculate an image at any desired TI. We refer to the collection of such images, computed at different TI values, as multi-TI images. Figures 4 and 5 illustrate this process. Figure 4 shows examples of the acquired MPRAGE and FGATIR images along with their estimated T1 and PD parameter maps. Using these maps, Figure 5 shows synthetic images computed at various TI values.
MPRAGE FGATIR T1111 Map PD Map
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Acquired MPRAGE and FGATIR images of one subject in our cohort of 23 subjects with their corresponding estimate of their T1 and PD parameter maps.
TI=600 TI=800 TI=1000 TI=1200
Ground Truth Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Example multi-TI images which we compute from the estimate of the parameter maps shown in Fig. 4.
2.5 Synthesis Model
To enable generation of multi-TI images when an FGATIR image is not acquired, we designed a synthesis network using a 2D U-Net [25] network to take commonly acquired MR images and synthesize T1 and PD maps, from which multi-TI images can be computed using Eq. 1. We trained the network by concatenating corresponding 2D slices from the MPRAGE, T2-w, and FLAIR images in any orientation into a three-channel input tensor. The outputs of this network are T1 and PD maps, which are available for training purposes from our cohort using the estimation step in Sec. 2.4. We use this 2D network to build a 3D volume by application of the network in all three cardinal orientations. We fuse these three outputs by taking their median, yielding 3D T1 and PD maps. In the encoder portion of our network, we use max pooling with a $2\times 2$ kernel and a stride of $2$ to reduce spatial dimensions. In the decoder, feature maps are upsampled and concatenated with corresponding encoded features, creating skip connections that help preserve spatial context at every downsampling resolution. Both the encoder and decoder employ 2D convolutions followed by instance normalization and LeakyReLU activation to stabilize training and introduce non-linearity. The network’s output is processed by a sigmoid function and then scaled by $4,000$, a modification we found to expedite convergence and allow parameter maps to reach realistic values more effectively than using a traditional ReLU activation. We optimized the network using the Adam optimizer with a learning rate of $\mathrm{10^{-6}}$.
[…]
Figure 6: Overview of the different training losses. Loss #1 is calculated by the summation of the L1 losses on the predicted and ground truth T1 and PD maps. Loss #2 is calculated by the summation of the L1 losses on the predicted and ground truth MPRAGE and FGATIR images.
2.6 Test time harmonization and imputation
Our synthetic T1 and PD maps are quantitative, making them sensitive to domain shifts in input images. Therefore, when applying our model to out-of-domain datasets or datasets with missing contrasts, we employ the HACA3 [20] MR harmonization algorithm to ensure that the input images resemble those used during training. HACA3 is an open source [^1] https://github.com/lianruizuo/haca3 unsupervised image harmonization approach for structural MR neuroimages. It does not require paired subject data for training. It uses an encoder-decoder structure to learn latent representations of anatomy, acquisition contrast, and image quality, enabling harmonization by combining these representations. HACA3 has the flexibility to handle multiple contrast MR images when harmonizing. It also has the ability to impute missing contrast MR images. When using HACA3, our target image contrasts are the T1-w MPRAGE, T2-w, and FLAIR images that we used to training our synthesis model. The encoded target images are available on our Github page: https://github.com/UponAcceptance. We use the synthetic images output as inputs to SyMTIC.
3 Experiments and Results
3.1 Comparison of training objectives for SyMTIC
First, we tested our models on the in-domain test set ($N=23$). An illustrative example of the predicted T1 and PD parameter maps, along with the corresponding synthetic MPRAGE and FGATIR images, is shown in Fig. 7 with the corresponding ground truth images (the acquired images and computed parameter maps). Table 1 reports the PSNR and SSIM for each image synthesized from each model. To ensure the evaluation focuses on meaningful image content, we excluded background voxels from the error calculations. Among the three models, we observed similar performances but Model #2 does give us the highest quantitative results for the four images. For statistical testing, we conducted a paired Wilcoxon test with Bonferroni correction. For the MPRAGE and T1 map, Model #2 PSNR and SSIM was significantly higher than Models #1 and #3. For the FGATIR, only the SSIM for Model #2 was significantly higher than Models #1 and #3. For the PD map, the SSIM for Model #2 was significantly higher than Model #1.
MPRAGE FGATIR T1 Map PD Map
Ground Truth Refer to caption Refer to caption Refer to caption Refer to caption
Model #1 Refer to caption Refer to caption Refer to caption Refer to caption
Model #2 Refer to caption Refer to caption Refer to caption Refer to caption
Model #3 Refer to caption Refer to caption Refer to caption Refer to caption
Figure 7: Ground truth images of a testing subject with the prediction images using the proposed SyMTIC models.
Table 1: Calculated PSNR and SSIM values on the synthetic images and parameter maps using the proposed SyMTIC models ($N=23$).
Model Loss #1 Loss #2 Image PSNR SSIM
Model #1 MPRAGE 40.33 $\pm$ 3.11 0.9512 $\pm$ 0.0133
FGATIR 26.91 $\pm$ 1.34 0.7334 $\pm$ 0.0472
T1 Map 22.10 $\pm$ 0.99 0.7804 $\pm$ 0.0247
PD Map 29.21 $\pm$ 0.84 0.8381 $\pm$ 0.0298
Model #2 MPRAGE 45.70 $\pm$ 5.67 0.9970 $\pm$ 0.0029
FGATIR 27.60 $\pm$ 2.27 0.7906 $\pm$ 0.0535
T1 Map 23.75 $\pm$ 1.73 0.8088 $\pm$ 0.0319
PD Map 29.82 $\pm$ 2.14 0.8516 $\pm$ 0.0365
Model #3 MPRAGE 42.35 $\pm$ 3.73 0.9564 $\pm$ 0.0127
FGATIR 26.99 $\pm$ 1.43 0.7408 $\pm$ 0.0492
T1 Map 22.31 $\pm$ 0.97 0.7906 $\pm$ 0.0228
PD Map 29.25 $\pm$ 1.39 0.8437 $\pm$ 0.0305
3.2 Multi-TI synthesis for in-domain testing dataset
Figure 8 shows examples of multi-TI images computed from the SyMTIC and ground truth T1 and PD maps shown in Fig. 7. Figure 9 provides a visual comparison of a acquired and synthetic MPRAGE and FGATIR images in the axial, sagittal, and coronal planes. We highlight DBS targets in these images, demonstrating that these structures are clearly visible in both the acquired and synthetic FGATIR images. This qualitative comparison supports the potential of our model to enhance DBS planning through synthetic FGATIR image generation.
TI=600 TI=800 TI=1000 TI=1200
Ground Truth Refer to caption Refer to caption Refer to caption Refer to caption
Model #1 Refer to caption Refer to caption Refer to caption Refer to caption
Model #2 Refer to caption Refer to caption Refer to caption Refer to caption
Model #3 Refer to caption Refer to caption Refer to caption Refer to caption
Figure 8: Synthetic multi-TI images of a test subject calculated using the predicted parameter maps from the SyMTIC models in Fig. 7.
Acquired MPRAGE
Synthetic MPRAGE
Acquired FGATIR
Synthetic FGATIR
Figure 9: Acquired MPRAGE, synthetic MPRAGE, acquired FGATIR, and synthetic FGATIR images (top to bottom). Arrows point to DBS targets that are less visible in MPRAGE images: blue and red for subthalamic nuclei, green for internal lamina.
3.3 Thalamus segmentation using THOMAS
It has been reported by Su et al., [7] that THOMAS achieves highest thalamus segmentation when using an FGATIR image. The goal of this experiment is to evaluate whether our synthetic FGATIR images could serve as a viable alternative to acquired FGATIR images for thalamus segmentation. By comparing THOMAS segmentation results across different image types, we aimed to assess whether our synthetic FGATIR images can bridge the gap between standard clinical imaging protocols and the enhanced contrast required for accurate thalamic delineation. This validation is particularly important for extending automated thalamic segmentation methods to datasets where FGATIR images are not available, ultimately broadening the clinical utility of standard MR imaging acquisitions. Since we do not have thalamus labels for our dataset, we used the THOMAS result on the acquired FGATIR image as the ground truth. We compared the THOMAS result on the acquired MPRAGE and FGATIR images from the three models with the THOMAS result on the acquired FGATIR. The Dice coefficients (DSC) are reported in Table 2. Our results indicate that THOMAS segments the thalamus more accurately using any of the synthetic FGATIR images than the acquired MPRAGE. This improvement is statistically significant across all models (Wilcoxon signed-rank test, p $<0.001$ for all comparisons).
Table 2: The entire thalamus DSC and mean DSC of the 12 classes for different input images to THOMAS.
Input Image DSC for Entire Thalamus Mean DSC of 12 Classes
Acquired MPRAGE 0.820 $\pm$ 0.059 0.708 $\pm$ 0.063
Synthetic FGATIR Model #1 0.894 $\pm$ 0.066 0.826 $\pm$ 0.064
Synthetic FGATIR Model #2 0.896 $\pm$ 0.067 0.828 $\pm$ 0.067
Synthetic FGATIR Model #3 0.878 $\pm$ 0.059 0.810 $\pm$ 0.096
3.4 Missing FLAIR image
We directly tested Models #1-3 on a dataset ($N=35$) that was acquired at the University of Maryland similarly to our training data, except the FLAIR image was either not acquired or of poor quality. We refer to this dataset as MMTI. This dataset also has MPRAGE and FGATIR images so it allowed us to evaluate our model performance on a large set of unseen data. To address the missing or low quality FLAIR image contrast, we used HACA3 to synthesize a FLAIR image using the acquired MPRAGE and T2-w images. An acquired MPRAGE image, acquired T2-w image, and synthetic FLAIR image from the MMTI dataset are shown in Fig. 10. HACA3 was used to impute the FLAIR image. From the predicted parameter maps, we calculated the multi-TI images shown in Fig. 11. This experiment shows that SyMTIC can handle datasets with missing FLAIR images by utilizing HACA3 for imputation.
MPRAGE T2-w FLAIR
Refer to caption Refer to caption Refer to caption
Figure 10: An acquired MPRAGE image, acquired T2-w image, and synthetic FLAIR image from the MMTI dataset. The synthetic FLAIR was imputed using HACA3 and the other two acquired images.
FGATIR TI=600 TI=800 TI=1000 TI=1200 MPRAGE
Ground Truth Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Model #1 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Model #2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Model #3 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 11: Calculated multi-TI images from the MMTI dataset using Models #1-3.
3.5 Synthesis for out-of-domain dataset
When our testing dataset was not acquired with a similar acquisition protocol to our UMD training dataset, we used HACA3 for image harmonization and imputation. An example of acquired and harmonized images from an out-of-domain dataset from people with multiple sclerosis (PwMS) are shown in Fig. 12. We used the harmonized images as input to Model #2. The predicted parameter maps and calculated multi-TI images for one subject are shown in Fig. 13. Since this dataset does not have paired T1-w images, we could not calculate ground truth parameter maps for comparison. However, we can assume our harmonized T1-w image to be an MPRAGE because HACA3 used an MPRAGE as the harmonization target. The PSNR and SSIM of the synthetic MPRAGE images in the PwMS dataset are $33.76\pm 2.28$ and $0.9967\pm 0.0012$, respectively. This experiment shows that by using HACA3 on out-of-domain data, SyMTIC is robust to changes in image acquisition parameters.
MPRAGE T2-w FLAIR
Acquired Refer to caption Refer to caption Refer to caption
Harmonized Refer to caption Refer to caption Refer to caption
Figure 12: Images acquired in the PwMS dataset for one subject (top row) that were not acquired using the same protocol as our UMD training dataset. The harmonized images (bottom row), which have a similar contrast to our training dataset, are the input to our model.
T1 Map PD Map MPRAGE FGATIR
Refer to caption Refer to caption Refer to caption Refer to caption
TI = 600ms TI = 800ms TI = 1000ms TI = 1200ms
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 13: Predicted parameter maps and multi-TI images for an out-of-domain testing set using Model #2.
4 Discussion and Conclusion
In this work, we proposed SyMTIC, a framework for multi-TI image synthesis using commonly acquired MR images: T1-w, T2-w, and FLAIR. Our results demonstrate that SyMTIC can accurately synthesize multi-TI images from clinically acquired MR images by estimating the T1 and PD parameter maps. The synthetic multi-TI images achieve accuracy comparable to those derived from acquired paired T1-w images, yielding improved differentiation of thalamic nuclei and other deep brain structures, particularly for TI values in the range of 400–800 ms. This finding aligns with prior studies that emphasize the benefits of short TI images for thalamic visualization and surgical planning for DBS [4].
Our evaluation of different training losses revealed that using Model #2 for the SyMTIC model provides the best balance between parameter estimation and image synthesis accuracy. We did not compare with direct synthesis using a U-Net because of its limited flexibility. Training a direct synthesis model would require us to only synthesize a fixed set of TI images. Our approach allows for customization of the TI value during inference. Future work may explore alternative loss functions that explicitly incorporate tissue-specific constraints, ensuring that synthesized images retain clinically meaningful tissue contrast.
One of the most significant outcomes of our study is the ability of SyMTIC to generalize to out-of-domain datasets with unknown acquisition parameters. Our results demonstrate that harmonization with HACA3 mitigates domain shift, enabling robust synthesis even when the input image contrast deviates from the training distribution. This robustness is particularly important for real-world clinical applications, where variations in scanner hardware, pulse sequences, and patient populations introduce inconsistencies that can degrade synthesis quality. Our findings suggest that incorporating harmonization techniques into synthetic MR imaging workflows can enhance their clinical utility by ensuring consistency across diverse datasets.
Additionally, our segmentation experiments using THOMAS indicate that synthetic FGATIR images provide superior thalamic segmentation accuracy compared to acquired MPRAGE images, with a mean DSC values of the entire thalamus improving from 0.820 (acquired MPRAGE) to 0.896 (synthetic FGATIR Model #2). This result is particularly important for applications such as DBS and MRgFUS, where accurate localization of deep gray matter structures is critical. While our study confirms the benefits of synthetic FGATIR images for thalamic segmentation, further validation with expert-labeled datasets would strengthen confidence in its clinical applicability.
Another key aspect of our approach is that we retain the skull region in our synthetic images rather than applying skull-stripping, which is common in many MR preprocessing pipelines. Although removal of the skull simplifies some analyses, retaining the skull provides a valuable anatomical context for applications where the interaction between the skull and brain is relevant [26]. For example, MRgFUS treatments rely on measurements of skull thickness and density to optimize ultrasound delivery, particularly in procedures for essential tremor and Parkinson’s disease. By preserving the skull in our synthetic images, SyMTIC ensures compatibility with studies investigating these effects, while still enabling downstream analyses of deep gray matter structures. Although we did not explicitly validate skull retention in our experiments, future work could explore its impact on MRgFUS targeting, skull-based registration techniques, and volumetric analyses in conditions affecting cranial morphology.
Our study also addresses the challenge of missing FLAIR MR images in clinical datasets. We demonstrate that SyMTIC remains effective as evidenced by our experiments with the MMTI dataset, where the synthetic FLAIR image imputed using HACA3 successfully enabled multi-TI synthesis. This capability is particularly relevant for retrospective studies, where incomplete imaging protocols limit the ability to perform advanced quantitative analysis. Future work could explore the integration of using HACA3 to impute other missing MR contrasts such as T1-w or T2-w.
While our framework focuses on T1 and PD parameter map synthesis, extending SyMTIC to estimate T2 maps would significantly broaden its applicability. The addition of T2 mapping would enable direct calculation of a wider range of imaging contrasts beyond the inversion recovery multi-TI images discussed in this work. Such an extension could further enhance applications in quantitative MRI, including relaxometry-based tissue characterization and lesion assessment in neurodegenerative diseases [12].
In summary, SyMTIC represents a novel approach for synthesizing multi-TI MR images from commonly acquired clinical sequences. Our method enables enhanced visualization of deep gray matter structures, improves segmentation accuracy, and generalizes well to datasets with varying acquisition protocols. By integrating harmonization and contrast imputation, SyMTIC offers a flexible solution for addressing the practical challenges of clinical imaging. Future work should focus on refining the synthesis process, expanding parameter estimation to include additional tissue properties, and validating the method across larger multi-site datasets to further establish its clinical utility.
5 Disclosures
The authors declare that there are no financial interests, commercial affiliations, or other potential conflicts of interest that could have influenced the objectivity of this research or the writing of this paper.
6 Code, Data, and Materials Availability
All code associated with the model development and testing is publicly available at https://github.com/UponAcceptance. The data utilized in this study were obtained from private datasets and not publicly available due to privacy concerns.
7 Acknowledgments
This work was partially supported by the National Science Foundation Graduate Research Fellowship under Grant No. DGE-2139757 (S. P. Hays) and National Cancer Institute (NCI) grants R01 CA253923 (L. Zuo) and R01 CA275015 (L. Zuo). It was also supported by the PCORI grant MS-1610-37115 (PIs: S. D. Newsome and E. M. Mowry). The statements in this publication are solely the responsibility of the authors and do not necessarily represent the views of PCORI, its Board of Governors or Methodology Committee.